OpenStudy (anonymous):

a population of inseects increase 10% each day,how many days will it take to double?

OpenStudy (anonymous):

solve \[(1.10)^t=2\] for t

OpenStudy (anonymous):

we can go right to the answer since \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] in your case \[t=\frac{\ln(2)}{\ln(1.1)}\] then a calculator

OpenStudy (anonymous):

two things you need to know for this problem first is that to increase a number by 10% you multiply it by \(1.1\) and second is how to solve when a variable is in the exponent. that was the second part